OpenStudy (anonymous):

Andrew believes the honor roll students at his school have an unfair advantage in being assigned to the math class they request. He asked 500 students at his school the following questions: "Are you on the honor roll?" and "Did you get the math class you requested?" The results are shown in the table below: Honor roll Not on honor roll Total Received math class requested 215 125 340 Did not get math class requested 80 80 160 Total 295 205 500 Help Andrew determine if all students at his school have an equal opportunity to get the math class they requested. @ganeshie8

OpenStudy (anonymous):

I did the problem and solved and got that 72.8 or 73% and 10 who recieved the math class and is on the honor roll and the 10% is for the students who are not on the honor but did recieve the class they wanted

OpenStudy (anonymous):

I got that mixed up those were for the students who were on the honor roll who did not get math class requested

ganeshie8 (ganeshie8):

P(Recieved math class) = ?

OpenStudy (anonymous):

Honor roll and Received math class requested = 215/295=72.8 or 73% Honor roll and Did not get math class requested= 80/295=10% Not on honor roll and Received math class requested 125/205=61% Not on honor roll and Not on honor roll = 80/205=39% @ganeshie8

OpenStudy (anonymous):

That is what I got.

ganeshie8 (ganeshie8):

thats right, what about the probability i have asked earlier ?

OpenStudy (anonymous):

um what do you mean?

ganeshie8 (ganeshie8):

whats the probability for `receiving math class` ? P("Receiving math class") = ?

OpenStudy (anonymous):

215

ganeshie8 (ganeshie8):

Look at the table, total how many students received math class ?

ganeshie8 (ganeshie8):

You need to use the last column...

OpenStudy (anonymous):

can you help me on another problem? @ganeshie8